Compute the steering angle (in degrees) required for an autonomous vehicle with pure pursuit lateral control for following the desired path based on the information below.

The lookahead distance is 15 m; the car length is 5 m; the angle between the vehicle’s body heading and the lookahead line is 60°.
1
Expert's answer
2020-12-07T03:46:56-0500

Look head distance=15 m="l_d" , length of car =L=5 m, the angle between the vehicle's body line and look head line is,"\alpha=60^o"

Now for steering angle we know that


"\delta=tan^{-1}(\frac{2Lsin \alpha}{l_d})=tan^{-1}(\frac{2\times 5\times sin 60}{15})=30^o"

So, here steering angle is 30o